# Local uniqueness and non-degeneracy of blowup solutions for regular Liouville systems

### v1

@@
$\Delta u_{i}=\sum_{j=1}^{n}a_{ij}\rho_{j}\Big{(}\frac{h_{j}e^{u_{j}}}{\int_{\Omega}h_{j}e^{u_{j}}}-1\Big{)}\mbox{ in }M\mbox{ for }i=1,\cdots,n,$
where the coefficient matrix $A=(a_{ij})_{n\times n}$ is nonnegative, $h_{1},\ldots,h_{n}$ are positive smooth functions, and $\rho_{1},\ldots,\rho_{n}$ are positive constants. For the blowup solutions, we establish their uniqueness and non-degeneracy based on natural assumptions.
@@ 1 Introduction and main results
$\Delta u_{i}=\sum_{j=1}^{n}a_{ij}\rho_{j}\Big{(}\frac{h_{j}e^{u_{j}}}{\int_{\Omega}h_{j}e^{u_{j}}}-1\Big{)}\mbox{ in }M\mbox{ for }i=1,\cdots,n.$ (1.1)
here $h_{1},...h_{n}$ are positive smooth functions, $\rho_{1},..,\rho_{n}$ are positive constants, $(a_{ij})_{n\times n}$ is a matrix satisfying
@@
    for $i=1,\cdots,n$, $\frac{h_{i}e^{u_{i}^{k}}}{\int_{\Omega}h_{i}e^{u_{i}^{k}}}\to\sum_{t=1}^{N}\alpha_{t}\delta_{p^{*}_{t}}$ as $k\to\infty$. Here, the constant $\alpha_{t}>0$ and $\delta_{q^{*}_{t}}$ is the Dirac measure supported by $\{q^{*}_{t}\}$ for $t=1,\cdots,N$.
@@
$M^{k}_{t}:=\max_{i\in I}\max_{x\in B_{\tau}(q_{t}^{*})}\Big{\{}u_{i}^{k}(x)-\ln\int_{\Omega}h_{i}e^{u_{i}^{k}(x)}dx\Big{\}}\quad\mbox{ and }\quad\varepsilon^{k}_{t}:=e^{-\frac{M^{k}_{t}}{2}}.$
@@ 3.1 An initial coordinate and classical estimates
$\max_{i}\sup_{|y|\leq{C/\varepsilon_{1}^{k}}}\frac{|w_{i}^{k}(y)|}{((\varepsilon_{1}^{k})^{\delta}\alpha^{k}+(\varepsilon_{1}^{k})^{m^{*}-2})(1+|y|)^{2\delta}}\Bigg{|}_{\partial B_{\tau/\varepsilon_{k}}}\to 0.$
@@
Here, we remark that $\xi_{j}^{k}\to V_{j}$ over any compact subset of $\mathbb{R}^{2}$ , which is due to the difference of the equations of $V_{1,i}^{k}$ and $\overline{V}_{2,i}^{k}$. Passing $k\to\infty$, we find that $\hat{w}_{i}^{k}\to cZ_{0,i}$ in $C_{loc}^{2}(\mathbb{R}^{2})$ for certain constant $c\in\mathbb{R}$. From the definition of $\eta_{k}$ we see that at least for one index $i_{1}$ we have $\hat{w}_{i_{1}}(0)=0$, thus $c=0$ and $\hat{w}_{i}=o(1)$ over any compact subset of $\mathbb{R}^{2}$. This rules out the possibility that $|r_{k}|<C$ for $C$ independent of $k$. But if $r_{k}\to\infty$, we use the Green’s representation of $\hat{w}^{k}_{i_{0}}(r_{k})$ to see that $\hat{w}^{k}_{i_{0}}(r_{k})=o(1)$, a contradiction to $\hat{w}^{k}_{i_{0}}(r_{k})=\pm 1$. The readers may read [38, pp.402-403] for more details in this argument. (3.1) is established. Claim 3.1 is established. $\Box$
@@ 3.2 Choosing a refined coordinate
In order to obtain a higher order expansion, we choose a suitable coordinate to make the boundary oscillation as small as possible. To do this, we proceed a reduction argument. Based on the results in the last subsection, we have the following preliminary estimate. First, we locate the point $p_{1,t}^{k}\in B_{\frac{\tau_{0}}{\varepsilon_{1}^{k}}}(0)$.
@@
$v_{i}^{1,k}(y)=\int_{\Omega_{1,k}}G_{k}(y,\eta)\Big{(}\sum_{j=1}^{n}a_{ij}H_{j}^{k}(\eta)e^{v_{i}^{1,k}(\eta)}\Big{)}d\eta-\int_{\partial\Omega_{1,k}}\partial_{\upsilon}G_{k}v_{i}^{1,k}d\mathcal{H}^{1}_{\eta}.$
@@
$\int_{\Omega_{1,k}}\frac{1}{2\pi}\ln\frac{|y|}{(\varepsilon_{1}^{k})^{-1}\tau_{1}^{k}}\Big{(}\sum_{j=1}^{n}a_{ij}H_{j}^{k}(\eta)e^{v_{i}^{1,k}(\eta)}\Big{)}d\eta+\int_{\Omega_{1,k}}\frac{1}{2\pi}\ln\frac{|y^{*}-\eta|}{|y-\eta|}\Big{(}\sum_{j=1}^{n}a_{ij}H_{j}^{k}(\eta)e^{v_{i}^{1,k}(\eta)}\Big{)}d\eta$
@@ 3.3 A refined expansion inside the bubbling disks
$\int_{\Omega_{M,k}}H_{i}^{M,k}(y)e^{v_{i}^{M,k}}=\int_{\Omega_{M,k}}H_{i}^{M,k}(0)e^{V_{j}^{M,k}}+T^{\prime}_{M^{\prime}}+O(\varepsilon^{(M^{\prime}+2)(m^{*}-2)})$
@@
$\int_{\Omega_{M,k}}H_{i}^{M,k}(y)e^{v_{i}^{M,k}}=\int_{\Omega_{M,k}}H_{i}^{M,k}(0)e^{V_{i}^{M,k}}+b_{i}^{k}(\varepsilon_{1}^{k})+T^{\prime}_{M^{\prime}}+O(\varepsilon^{(M^{\prime}+2)(m^{*}-2)})$
@@
$T^{\prime}_{0,M^{\prime}},\,T^{\prime}_{M^{\prime}}=o_{\tau}(1)\varepsilon^{m^{*}-2}$ and is $C^{2}$ depends on $p_{i}^{M,k}$, $h_{i}$ and its derivatives up to $M^{\prime}+3$ order. The functions $H_{i}^{M,k}$ and $h_{i}^{M,k}$ are defined as in Subsection 3.2.
@@
$=\Delta(V_{i}^{M,k}+c_{1,1,i,k})+\sum_{j}a_{ij}H_{j}^{M,k}(0)e^{V_{j}^{M,k}+c_{1,1,j,k}+\ln\frac{H_{j}^{M,k}(\varepsilon_{1}^{k}y)}{H_{j}^{M,k}(0)}}$
$=\Delta(V_{i}^{M,k}+c_{1,1,i,k})+\sum_{j}a_{ij}H_{j}^{M,k}(0)e^{V_{j}^{M,k}}\Bigg{(}1+c_{1,1,j,k}+\ln\frac{H_{j}^{M,k}(\varepsilon_{1}^{k}y)}{H_{j}^{M,k}(0)}$
$\quad+\frac{1}{2}\Bigg{(}c_{1,1,j,k}+\ln\frac{H_{j}^{M,k}(\varepsilon_{1}^{k}y)}{H_{j}^{M,k}(0)}\Bigg{)}^{2}+O((\varepsilon_{1}^{k})^{3}|y|^{3})\Bigg{)}.$ (3.32)
@@
$(\varepsilon_{1}^{k})^{2}\int_{\Omega_{M,k}}e^{V_{i}^{M,k}}Z_{0,i}|y|^{2}dy=4\pi(\varepsilon_{1}^{k})^{2}\int_{0}^{\tau_{k}^{M}/\varepsilon_{1}^{k}}e^{V_{j}}r^{3}dr=O\Big{(}\varepsilon^{2}\ln\frac{1}{\varepsilon}\Big{)}.$ (3.37)
@@
Next we obtain a $M^{\prime}$-th order approximation recursively. Denoting this term by $c_{M^{\prime},i}^{k}$. It is known that
$|c_{M^{\prime},i}^{k}(y)|=O((\varepsilon_{1}^{k})^{M^{\prime}}(1+|y|)^{M^{\prime}{+\delta}})\,\mbox{ in }\,\Omega_{M,k}$ (3.38)
@@
$\Delta(V_{i}^{M,k}+c_{M^{\prime},i}^{k})+\sum_{j}a_{ij}H_{j}^{M,k}(y)e^{V_{i}^{M,k}+c_{M^{\prime},i}^{k}}=:A_{M^{\prime},i}+B_{M^{\prime},i}$
@@
$A_{M^{\prime},i}(y)=O((\varepsilon_{1}^{k})^{M^{\prime}+1}|y|^{M^{\prime}+1}(1+|y|)^{-m^{*}{+\delta}})\,\,\mbox{and}\,\,B_{M^{\prime},i}(y)=O((\varepsilon_{1}^{k})^{M^{\prime}}(1+|y|)^{3M^{\prime}-(M^{\prime}+1)m^{*}{+\delta}})$
@@
$\left\{\begin{array}[]{lr}\Delta W^{*}_{i,k}+\sum_{j}a_{ij}H_{j}^{M,k}(y)e^{V_{i}^{M,k}}W^{*}_{j,k}=O((\varepsilon_{1}^{k})^{M^{\prime}+1}(1+|y|)^{3M^{\prime}-(1+M^{\prime})m^{*}{+\delta}})\mbox{ in }\Omega_{M,k},\\ W^{*}_{i,k}|_{\partial\Omega_{M,k}}=\mbox{constant}\mbox{ for }i=1,\cdots,n.\end{array}\right$. (3.42)
Denoting $(c_{M,i}+\phi_{1,i}^{**}-\phi_{2,i}^{**})_{0}-c_{2,0,i,k}=c_{*,0,i,k}$. Here, $(u)_{0}$ represents the radial part of $u$. A routine estimate implies that
$|W_{i,k}^{*}(y)|=O((\varepsilon_{1}^{k})^{M^{\prime}+1}(1+|y|)^{2+3M^{\prime}-(1+M^{\prime})m^{*}{+\delta}})$
@@
$\int_{\Omega_{M,k}}H_{j}^{M,k}(y)e^{v_{j}^{M,k}}$ $=\int_{\Omega_{M,k}}H_{j}^{M,k}(0)e^{V_{j}^{M,k}}+\int_{\partial\Omega_{M,k}}\frac{\partial c_{2,0,i,k}}{\partial{\bf n}}+\int_{\partial\Omega_{M,k}}\frac{\partial c_{*,0,i,k}}{\partial{\bf n}}+O(\varepsilon_{k}^{(M^{\prime}+2)(m^{*}-2){-\delta}})$
$=\int_{\Omega_{M,k}}H_{j}^{M,k}(0)e^{V_{j}^{M,k}}+T^{\prime}_{0,M^{\prime}}+O(\varepsilon_{k}^{(M^{\prime}+2)(m^{*}-2){-\delta}}).$
Here, the term $T^{\prime}_{0,M^{\prime}},\,T^{\prime}_{M^{\prime}}=o_{\tau}(1)\varepsilon^{m^{*}-2}$ and is $C^{2}$ depends on $p_{i}^{M,k}$, $h_{i}$ and its derivatives up to $M^{\prime}+3$ order. As similar argument can be proceed for the case $m^{*}=4$. The proof is completed.
@@ 3.4 A refined estimate outside the bubbling disks
$u_{i}^{k}(x)$ $=\Bigg{(}1-\frac{m_{i}^{k}}{2}\Bigg{)}M_{k,t}-\ln(\rho_{i}^{k}h_{i}^{k}(q_{t}^{k}))-2\pi m_{i}^{k}G^{*}(q_{t}^{M,k};q_{t}^{M,k})+I_{t}^{k}+T_{M}^{**}+O(\varepsilon^{M(m^{*}-2){-\delta}})$
@@
$\int_{M}u_{i}^{k}=-\frac{m_{itv}^{k}-2}{2}\ln\frac{1}{\varepsilon_{t}^{k}}-2\pi m_{itv}^{k}G^{*}(q_{t}^{M,k};q_{t}^{M,k})+I_{t}^{k}-\ln(\rho_{i}^{k}h_{i}^{k}(q_{t}^{M,k}))+T_{M}^{\mbox{Avg},k}+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
Here, $T_{M}^{**}$ depends $C^{2}$-continuously in $p_{t}^{M,k}$, $\varepsilon_{t}$ and the derivatives of $h_{i}^{k}$ up to $M+3$-th order.
@@
$=:m_{itv}^{k}G(x,q_{t}^{M,k})+T_{M}^{\mbox{out},k}+O(\varepsilon^{(M^{\prime}+2)(m^{*}-2){-\delta}}).$ (3.46)
Here, $T_{M}^{\mbox{out},k}$ is a term of order $o_{\tau}(\varepsilon^{m^{*}-2})$. Notice that we used the frequency analysis in Subsection 3.3. Now we derive the expansion outside the bubbling disks. By a classical estimate [26, Proposition 3.1], we get
@@
$u_{it}^{k}(x)=-m_{itv}^{k}\ln|x|-(m_{itv}^{k}-2)\ln\frac{1}{\varepsilon_{t}^{k}}+I_{t}^{k}+T^{\mbox{in},k}_{M^{\prime}}-\ln(\rho_{i}^{k}h_{i}^{k}(q_{t}^{M,k}))+O(\varepsilon^{(M^{\prime}+2)(m^{*}-2){-\delta}}).$ (3.48)
Matching (3.47) with (3.48), we get
@@
$\Bigg{(}\frac{\varepsilon^{k}_{t}}{\varepsilon^{s}_{k}}\Bigg{)}^{m_{i}^{k}-2}=\frac{e^{H_{i,t}(q^{M,k})}}{e^{H_{i,s}(q^{M,k})}}+T_{i,M,t,s}^{OUT}+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
Here, $T_{i,M,t,s}^{OUT}$ depends $C^{2}$-continuously in $p_{t}^{M,k}$, $\varepsilon_{t}$ and the derivatives of $h_{i}^{k}$ up to $M+3$-th order for $t=1,\cdots,N$.
@@
$\tau_{k}^{M}=\tau_{0}+T^{\tau}_{M,t,k}+O(\varepsilon^{M(m^{*}-2)-\delta}).$
Here, $T_{M,t,k}^{\tau}=O(\varepsilon^{2})$ and depends $C^{2}$-continuously in $p_{t}^{M,k}$, $\varepsilon_{t}$ and the derivatives of $h_{i}^{k}$ up to $M+3$-th order for $t=1,\cdots,N$.

### v2

@@
$\Delta u_{i}=\sum_{j=1}^{n}a_{ij}\rho_{j}\Big(\frac{h_{j}e^{u_{j}}}{\int_{M}h_{j}e^{u_{j}}}-1\Big)\mbox{ in }M\mbox{ for }i=1,\cdots,n,$
where $M$ is a compact Riemann surface, the coefficient matrix $A=(a_{ij})_{n\times n}$ is nonnegative, $h_{1},\ldots,h_{n}$ are positive smooth functions, and $\rho_{1},\ldots,\rho_{n}$ are positive constants. For the blowup solutions, we establish their uniqueness and non-degeneracy based on natural assumptions.
@@ 1 Introduction and main results
$\Delta u_{i}=\sum_{j=1}^{n}a_{ij}\rho_{j}\Big(\frac{h_{j}e^{u_{j}}}{\int_{M}h_{j}e^{u_{j}}}-1\Big)\mbox{ in }M\mbox{ for }i=1,\cdots,n.$ (1.1)
Here, $M$ is a compact Riemann surface, $h_{1},...h_{n}$ are positive smooth functions, $\rho_{1},..,\rho_{n}$ are positive constants, $(a_{ij})_{n\times n}$ is a matrix satisfying
@@
    for $i=1,\cdots,n$, $\frac{h_{i}e^{u_{i}^{k}}}{\int_{\Omega}h_{i}e^{u_{i}^{k}}}\to\sum_{t=1}^{N}\alpha_{t}\delta_{q^{*}_{t}}$ as $k\to\infty$. Here, the constant $\alpha_{t}>0$ and $\delta_{q^{*}_{t}}$ is the Dirac measure supported by $\{q^{*}_{t}\}$ for $t=1,\cdots,N$.
@@
$M^{k}_{t}:=\max_{i\in I}\max_{x\in B_{\tau}(q_{t}^{*})}\Big\{u_{i}^{k}(x)-\ln\int_{M}h_{i}e^{u_{i}^{k}(x)}dx\Big\}\quad\mbox{ and }\quad\varepsilon^{k}_{t}:=e^{-\frac{M^{k}_{t}}{2}}.$
@@ 2 Preliminary Results
  • We use $M$ to denote a generically large integer which might be different from line to line;
  • We frequently use higher-order expansions and their continuous dependence on parameters, while the specific expression of the expansion itself is irrelevant. Therefore in this paper, we always denote the sum of higher-order terms as $T_{M,k}$ with certain superscripts or subscripts. We refer, for instance, Theorems 3.1 and 3.1, Propositions 3.3 and 3.4 and Corollaries 3.1 and 3.2.
@@ 3.1 An initial coordinate and classical estimates
$\max_{i}\frac{|w_{i}^{k}(y)|}{((\varepsilon_{1}^{k})^{\delta}\alpha^{k}+(\varepsilon_{1}^{k})^{m^{*}-2})(1+|y|)^{2\delta}}\Bigg|_{\partial B_{\tau/\varepsilon_{k}}}\to 0.$
@@
Here, we remark that $\xi_{j}^{k}\to V_{j}$ over any compact subset of $\mathbb{R}^{2}$ , which is due to the difference of the equations of $V_{1,i}^{k}$ and $\overline{V}_{2,i}^{k}$. Passing $k\to\infty$, we find that $\hat{w}_{i}^{k}\to cZ_{0,i}$ in $C_{loc}^{2}(\mathbb{R}^{2})$ for certain constant $c\in\mathbb{R}$. From the definition of $\eta_{k}$ we see that at least for one index $i_{1}$ we have $\hat{w}_{i_{1}}(0)=0$, thus $c=0$ and $\hat{w}_{i}=o(1)$ over any compact subset of $\mathbb{R}^{2}$. This rules out the possibility that $|r_{k}|<C$ for $C$ independent of $k$. But if $r_{k}\to\infty$, we use the Green’s representation of $\hat{w}^{k}_{i_{0}}(r_{k})$ to see that $\hat{w}^{k}_{i_{0}}(r_{k})=o(1)$, a contradiction to $\hat{w}^{k}_{i_{0}}(r_{k})=\pm 1$. The readers may read [38, pp.402-403] for more details in this argument. (3.1) is established. Claim 3.1 is established.
$\Box$
@@ 3.2 Choosing a refined coordinate
In order to obtain a higher order expansion, we choose a suitable coordinate to make the boundary oscillation as small as possible. To do this, we proceed a reduction argument. Based on the results in the last subsection, we have the following preliminary estimate. First, we locate the point $p_{1,t}^{k}\in B_{\tau_{0}/\varepsilon_{1}^{k}}(0)$.
@@
$v_{i}^{1,k}(y)=\int_{\Omega_{1,k}}G_{k}(y,\eta)\Big(\sum_{j=1}^{n}a_{ij}H_{j}^{1,k}(\eta)e^{v_{i}^{1,k}(\eta)}\Big)d\eta-\int_{\partial\Omega_{1,k}}\partial_{\upsilon}G_{k}v_{i}^{1,k}d\mathcal{H}^{1}_{\eta}.$
@@
$\int_{\Omega_{1,k}}\frac{1}{2\pi}\ln\frac{|y|}{(\varepsilon_{1}^{k})^{-1}\tau_{1}^{k}}\Big(\sum_{j=1}^{n}a_{ij}H_{j}^{1,k}(\eta)e^{v_{i}^{1,k}(\eta)}\Big)d\eta+\int_{\Omega_{1,k}}\frac{1}{2\pi}\ln\frac{|y^{*}-\eta|}{|y-\eta|}\Big(\sum_{j=1}^{n}a_{ij}H_{j}^{1,k}(\eta)e^{v_{i}^{1,k}(\eta)}\Big)d\eta$
@@ 3.3 A refined expansion inside the bubbling disks
$\int_{\Omega_{M,k}}H_{i}^{M,k}(y)e^{v_{i}^{M,k}}=\int_{\Omega_{M,k}}H_{i}^{M,k}(0)e^{V_{j}^{M,k}}+T^{in}_{M^{\prime},k}+O(\varepsilon^{(M+2)(m^{*}-2)})$
@@
$\int_{\Omega_{M,k}}H_{i}^{M,k}(y)e^{v_{i}^{M,k}}=\int_{\Omega_{M,k}}H_{i}^{M,k}(0)e^{V_{i}^{M,k}}+b_{i}^{k}(\varepsilon_{1}^{k})+T^{in,*}_{M,k}+O(\varepsilon^{(M^{\prime}+2)(m^{*}-2)}).$
@@
$T^{in}_{M,k},T^{in,*}_{M,k}=o_{\tau}(1)\varepsilon^{m^{*}-2}$ and is $C^{2}$ depends on $p_{i}^{M,k}$, $h_{i}$ and its derivatives up to $M+3$ order. The functions $H_{i}^{M,k}$ and $h_{i}^{M,k}$ are defined as in Subsection 3.2.
@@
$=\Delta(V_{i}^{M,k}+c_{1,1,i,k})+\sum_{j}a_{ij}H_{j}^{M,k}(0)e^{V_{j}^{M,k}+c_{1,1,j,k}+\ln\frac{H_{j}^{M,k}(y)}{H_{j}^{M,k}(0)}}$
$=\Delta(V_{i}^{M,k}+c_{1,1,i,k})+\sum_{j}a_{ij}H_{j}^{M,k}(0)e^{V_{j}^{M,k}}\Bigg(1+c_{1,1,j,k}+\ln\frac{H_{j}^{M,k}(y)}{H_{j}^{M,k}(0)}$
$\quad+\frac{1}{2}\Bigg(c_{1,1,j,k}+\ln\frac{H_{j}^{M,k}(y)}{H_{j}^{M,k}(0)}\Bigg)^{2}+O((\varepsilon_{1}^{k})^{3}|y|^{3})\Bigg).$ (3.32)
@@
$(\varepsilon_{1}^{k})^{2}\int_{\Omega_{M,k}}e^{V_{i}^{M,k}}Z_{0,i}|y|^{2}dy\leq 4\pi(\varepsilon_{1}^{k})^{2}\int_{0}^{\tau_{k}^{M}/\varepsilon_{1}^{k}}e^{V_{j}}r^{3}dr=O\Big(\varepsilon^{2}\ln\frac{1}{\varepsilon}\Big).$ (3.37)
@@
Next we obtain a $M$-th order approximation recursively. Denoting this term by $c_{M,i}^{k}$. It is known that
$|c_{M,i}^{k}(y)|=O((\varepsilon_{1}^{k})^{M}(1+|y|)^{M+\delta})\,\mbox{ in }\,\Omega_{M,k}$ (3.38)
@@
$\Delta(V_{i}^{M,k}+c_{M,i}^{k})+\sum_{j}a_{ij}H_{j}^{M,k}(y)e^{V_{i}^{M,k}+c_{M,i}^{k}}=:A_{M,i}+B_{M,i}$
@@
$A_{M,i}(y)=O((\varepsilon_{1}^{k})^{M+1}|y|^{M^{\prime}+1}(1+|y|)^{-m^{*}{+\delta}})\,\,\mbox{and}\,\,B_{M,i}(y)=O((\varepsilon_{1}^{k})^{M}(1+|y|)^{3M-(M+1)m^{*}{+\delta}})$
@@
$\left\{\begin{array}[]{lr}\Delta W^{*}_{i,k}+\sum_{j}a_{ij}H_{j}^{M,k}(y)e^{\overline{V_{i}^{M,k}}}W^{*}_{j,k}=O((\varepsilon_{1}^{k})^{M+1}(1+|y|)^{3M-(1+M)m^{*}{+\delta}})\mbox{ in }\Omega_{M,k},\\ W^{*}_{i,k}|_{\partial\Omega_{M,k}}=\mbox{constant}\mbox{ for }i=1,\cdots,n.\end{array}\right$. (3.42)
Here, $\overline{V_{i}^{M,k}}$ is due to Mean Value Theorem. Denoting $(c_{M,i}+\phi_{1,i}^{**}-\phi_{2,i}^{**})_{0}-c_{2,0,i,k}=c_{*,0,i,k}$. Here, $(u)_{0}$ represents the radial part of $u$. A routine estimate implies that
$|W_{i,k}^{*}(y)|=O((\varepsilon_{1}^{k})^{M+1}(1+|y|)^{2+3M-(1+M)m^{*}{+\delta}})$
@@
$\int_{\Omega_{M,k}}H_{j}^{M,k}(y)e^{v_{j}^{M,k}}$ $=\int_{\Omega_{M,k}}H_{j}^{M,k}(0)e^{V_{j}^{M,k}}+\int_{\partial\Omega_{M,k}}\frac{\partial c_{2,0,i,k}}{\partial{\bf n}}+\int_{\partial\Omega_{M,k}}\frac{\partial c_{*,0,i,k}}{\partial{\bf n}}+O(\varepsilon_{k}^{(M+2)(m^{*}-2){-\delta}})$
$=\int_{\Omega_{M,k}}H_{j}^{M,k}(0)e^{V_{j}^{M,k}}+T_{M,k}^{in}+O(\varepsilon_{k}^{(M+2)(m^{*}-2){-\delta}}).$
Here, the term $T^{in}_{M,k}=o_{\tau}(1)\varepsilon^{m^{*}-2}$ and is $C^{2}$ depends on $p_{i}^{M,k}$, $h_{i}$ and its derivatives up to $M^{\prime}+3$ order. As similar argument can be proceed for the case $m^{*}=4$. The proof is completed.
@@ 3.4 A refined estimate outside the bubbling disks
$u_{i}^{k}(x)$ $=\Bigg(1-\frac{m_{i}^{k}}{2}\Bigg)M_{k,t}-\ln(\rho_{i}^{k}h_{i}^{k}(q_{t}^{k}))-2\pi m_{i}^{k}G^{*}(q_{t}^{M,k};q_{t}^{M,k})+I_{t}^{k}+T_{M,k}^{out}+O(\varepsilon^{M(m^{*}-2){-\delta}})$
@@
$\int_{M}u_{i}^{k}=-\frac{m_{itv}^{k}-2}{2}\ln\frac{1}{\varepsilon_{t}^{k}}-2\pi m_{itv}^{k}G^{*}(q_{t}^{M,k};q_{t}^{M,k})+I_{t}^{k}-\ln(\rho_{i}^{k}h_{i}^{k}(q_{t}^{M,k}))+T_{M,k}^{avg}+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
Here, $T_{M,k}^{out}$ and $T_{M,k}^{avg}$ depend $C^{2}$-continuously in $p_{t}^{M,k}$, $\varepsilon_{t}$ and the derivatives of $h_{i}^{k}$ up to $M+3$-th order.
@@
$=:m_{itv}^{k}G(x,q_{t}^{M,k})+T_{M,k}^{G}+O(\varepsilon^{(M^{\prime}+2)(m^{*}-2){-\delta}}).$ (3.46)
Here, $T_{M,k}^{G}$ is a term of order $o_{\tau}(\varepsilon^{m^{*}-2})$. Notice that we used the frequency analysis in Subsection 3.3. Now we derive the expansion outside the bubbling disks. By a classical estimate [26, Proposition 3.1], we get
@@
$u_{it}^{k}(x)=-m_{itv}^{k}\ln|x|-(m_{itv}^{k}-2)\ln\frac{1}{\varepsilon_{t}^{k}}+I_{t}^{k}+T^{u,in}_{M^{\prime},k}-\ln(\rho_{i}^{k}h_{i}^{k}(q_{t}^{M,k}))+O(\varepsilon^{(M^{\prime}+2)(m^{*}-2){-\delta}}).$ (3.48)
Here, $T^{u,in}_{M^{\prime},k}$ is a higher order expansion of $u_{it}^{k}$. Matching (3.47) with (3.48), we get
@@
$\Bigg(\frac{\varepsilon^{k}_{t}}{\varepsilon^{s}_{k}}\Bigg)^{m_{i}^{k}-2}=\frac{e^{H_{i,t}(q^{M,k})}}{e^{H_{i,s}(q^{M,k})}}+T_{M,k,i,t,s}^{ratio}+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
Here, $T_{M,k,i,t,s}^{ratio}$ depends $C^{2}$-continuously in $p_{t}^{M,k}$, $\varepsilon_{t}$ and the derivatives of $h_{i}^{k}$ up to $M+3$-th order for $t=1,\cdots,N$.
@@
$\tau_{k}^{M}=\tau_{0}+T^{\tau}_{M,k,t}+O(\varepsilon^{M(m^{*}-2)-\delta}).$
Here, $T_{M,k,t}^{\tau}=O(\varepsilon^{2})$ and depends $C^{2}$-continuously in $p_{t}^{M,k}$, $\varepsilon_{t}$ and the derivatives of $h_{i}^{k}$ up to $M+3$-th order for $t=1,\cdots,N$.